OpenStudy (anonymous):

2/3x=12 Help again .-. pleaseee

jimthompson5910 (jim_thompson5910):

the equation is \[\large \frac{2}{3}x = 12\] right?

OpenStudy (anonymous):

Righhht

jimthompson5910 (jim_thompson5910):

first multiply both sides by 3 to get \[\large \frac{2}{3}x = 12\] \[\large 3*\frac{2}{3}x = 3*12\] \[\large 2x = 36\] what's next?